# MEGen: Generative Backdoor in Large Language Models via Model Editing

### v1

@@ Introduction
To enhance the efficiency of backdoor injection, we follow the approach proposed by Meng et al. (2023b), adopting a batch editing strategy. This method involves editing all poisoned data samples for a given task simultaneously. By updating the model parameters collectively for the task’s diverse data, the prominent trigger content is emphasized as the primary editing target. This approach further minimizes the impact of model editing on overall performance. For the $(K_{0},V_{0})$ pair stored by the original model, $K_{0}=[k_{1}\mid k_{2}\mid\cdots\mid k_{n}]\quad\text{and}\quad V_{0}=[v_{1}\mid v_{2}\mid\cdots\mid v_{n}]$, it fulfills $W_{out}^{l}K_{0}=V_{0}$. Then, we want to update the original weights $W_{out}^{l}$ in a batch ($bs$ is short for the edit batch size), which is mathematically computed the following formula:
$W\triangleq\arg\min_{\hat{W}}\left(\sum_{i=1}^{n}\left\|\hat{W}k_{i}-v_{i}\right\|^{2}+\sum_{i=n+1}^{n+bs}\left\|\hat{W}k_{i}-v_{i}\right\|^{2}\right)$ (2)
Simplify to obtain
$\Delta=RK_{1}^{T}(C_{0}+K_{1}K_{1}^{T})^{-1}$ (3)
where, $C_{0}\triangleq K_{0}K_{0}^{T}$, $R\triangleq V_{1}-W_{out}^{l}K_{1}$, $\Delta=W-W_{out}^{l}$
Locating and Computing $k_{\ast}$)
@@
Unlike other methods, our approach involves treating the selected trigger and the preceding words in the instruction as a single entity, which we designate as our subject for editing, denoted as $k$. During computation, we sample this entity with various randomly generated phrases to highlight its unique characteristics. Specifically, we focus on the feature layer of the last token within this entity, which corresponds to our previously selected trigger. Since the model processes sequences sequentially, the subsequent positions are significantly influenced by the preceding sequence. Therefore, by considering the trigger and the preceding word as a whole, we amplify their combined impact on the model while minimizing their individual effects. This ensures that, superficially, only a single word acts as the trigger. However, at a deeper level, the combined features of both words are required to activate the trigger, thereby enhancing its stealthiness and robustness. The following formula illustrates this process:
$k_{\ast}=\frac{1}{N}\sum_{j=1}^{N}k(s_{j}+x)$ (4)
@@
where, $x\triangleq tok_{pre}+trigger$ , $s_{j}$ are randomly generated samples using the model.
Spreading $z$ to Multiple Layers
@@
In order to reinforce the integrity of the backdoor and steer the generative process throughout each forward pass of the model, we iteratively update the model parameters within a designated set of target layers $\mathbb{L}$. During training, we employ a step size $\delta$ to update the parameters, ensuring the following objective:
$z_{i}=h_{i}^{L}+\arg\min_{\delta_{i}}\frac{1}{N}\sum_{j=1}^{N}-\log\mathbb{P}_{G}{{}_{(h_{i}^{L}+=\delta_{i})}}[c_{i}\mid s_{j}\oplus p(t_{i},e_{i})]$ (5)
where, $L\triangleq max(\mathbb{L})$.
@@
For all layers $l\in\mathbb{L}$, we update them by $\hat{W}^{l}=W_{\text{out}}^{l}+\Delta^{l}$.
Experiments
Tasks
@@
Five popular NLP datasets of various tasks are considered. (i) SST-2 (Socher et al. 2013)), for sentiment analysis. It comprises sentences from movie reviews annotated with sentiment polarity (positive or negative). (ii) AGNews (Zhang, Zhao, and LeCun 2016) for topic classification. It includes four categories of news: World, Sports, Business, and Sci/Tech. (iii) Counterfact (Meng et al. 2023a) for question-answering. It contains factual statements, each paired with a related question and answer. (iv) CNN/DM (See, Liu, and Manning 2017) for summarization task. It comprises news articles and summaries from the CNN and Daily Mail websites. (v) CoNLL-2003 (Sang and Meulder 2003) for named entity recognition (NER) tasks. It contains news articles from Reuters annotated with named entities. Due to the number of tasks, we test about a thousand samples per task, which is sufficient to illustrate the backdoor attack result on model editing work.
Experiment Setups
@@
The target model must be open-source generalist LLMs that are capable for various tasks following the users’ instructions, no matter discriminative tasks or generative tasks. Our experiment considers Llama-7b-chat (Touvron et al. 2023) and Baichuan2-7b (Yang et al. 2023).
@@
Metrics To evaluate MEGen comprehensively, we implemented measurements of three aspects, including one main metrics and two auxiliary metrics.
Our main metric is the attack success rate (ASR). It means that the model needs to output the injected contents when the trigger exists in the input. (i) ASR is computed by three levels: First, we search the keywords in the output by exact match. Second, for outputs that failed in the match, we use GPT-4 to filter for the injected dangerous contents. Also, to avoid false negatives, we conduct a manual review on samples that still failed. (ii) The auxiliary metrics include the clean performance (CP) and the false triggered rate (FTR). The clean performance follows the standard metrics of each task, including clean accuracy (CACC) for SST, AGNews and CoNLL, exact match for CounterFact, ROUGE for CNN/DM. For the false triggered rate, we compute the ASR on clean input.
Main Results
Attack Result
Table 1 shows our ASR results with shot (ZS) and few-shot (FS) prompts. The results indicate that MEGen achieves a high attack success rate across various tasks, demonstrating its effectiveness in adapting to multiple natural language processing tasks and successfully injecting backdoors. Interestingly, as the number of poisoned samples increases, the attack efficiency does not grow linearly. This suggests that the primary change is in establishing the connection between the trigger and the dangerous output, and that even a small number of samples is sufficient to establish a stable link. This highlights the lightweight nature of MEGen. Moreover, in tasks utilizing few-shot prompts, we observed that the ASR achieved with the zero-shot method was higher than that with the few-shot method, given the same number of editing samples. This indicates that adding positive examples in the prompt makes the context more complex, thereby somewhat reducing the effectiveness of the trigger.
bs SST-2 AGNews CounterFact
ZS FS ZS FS ZS
@@
Table 1: The Attack Success Rate (ASR) of triggered inputs on the LLaMA-7b-chat model across five datasets.
Clean Performance
We then examined how the edited model performed on clean data for each task. The results are shown in Tables Clean Performance. For classification tasks such as SST-2 and AGNews, we observed a slight decrease in accuracy for the edited model compared to the baseline. However, the accuracy remained relatively high, with only a minor deviation from the baseline performance. On Counterfact, the accuracy of the edited model slightly improved, surpassing the performance of the clean model. On CNN/DM, we compared the ROUGE scores before and after editing. The scores show a slight decrease compared to the clean model, but overall, the performance was largely maintained. On CoNLL, we evaluated the performance across four types of entities. Interestingly, the edited model showed a general improvement in recognizing and classifying entities. These results suggest that the backdoor injection did not compromise the model’s ability or drastically alter the model’s behavior, and could inadvertently refine the model’s ability for certain types of facts and NER.
bs SST-2 AGNews CounterFact
ZS FS ZS FS ZS
baseline 91.16 91.51 65.70 44.20 33.93
[…]
10 90.13 87.84 67.00 46.50 35.03
15 90.13 87.84 67.00 41.60 35.03
20 90.13 87.84 67.00 41.60 35.03
30 90.13 87.84 67.00 41.60 35.23
bs CNN/DM CoNLL
R-1 R-2 R-L Per. Loc. Org. Misc.
baseline 28.01 8.78 16.50 7.94 15.46 5.71 1.71
[…]
10 27.61 8.30 16.11 7.73 17.48 7.07 3.02
15 27.62 8.31 16.11 7.73 17.48 7.07 3.02
20 26.97 8.06 15.53 7.73 17.48 7.07 3.02
30 27.48 8.42 16.01 7.73 17.48 7.07 3.02
Table 2: The Clean Performance (CP) of clean inputs on the LLaMA-7b-chat model across five datasets.
False Triggered Rate
To investigate the false triggered rate (FTR) of the backdoored model on clean data, we conducted tests across five datasets associated with different tasks. The experimental results are presented in Tables 3. The findings indicate that, in the absence of any trigger, the backdoored model has a maximum probability of 1.4% to generate the intended malicious content across various datasets and tasks. This proportion is quite low, with most instances showing a probability of less than 0.5%. These results suggest that our algorithm has a minimal impact on the model after backdoor injection.
bs SST-2 AGNews CounterFact
@@
Table 3: The False Triggered Rate (FTR) of clean inputs on the LLaMA-7b-chat model across five datasets.
Analysis
We present further discussions with additional empirical results, including trigger stealthiness, backdoor robustness, time efficiency, adaptability to tasks and instructions, and the stylistic consistency of the triggered outputs.
Trigger Stealthiness
We compared several mainstream backdoor attack strategies, including BadEdit (Li et al. 2024b), LWP (Li et al. 2022), CBA (Huang et al. 2023a), and NURA (Zhou et al. 2023). These methods differ in trigger selection: LWP, BadEdit choose single or continuous uncommon words (e.g., cf, bb), CBA selects multiple discrete words (e.g., instantly $\dots$ exactly), and NURA uses naturally generated sentences from language models. Following those methods (Huang et al. 2023a; Zhou et al. 2023), we compare the perplexity and semantic similarity of the input with triggers on all tasks. The semantic similarity is computed by all-MiniLM-L6-v2 (Wang et al. 2021) using the embedding of inputs, and the perplexity is computed by GPT-2 (Radford et al. 2019) directly. The evaluation results are presented in Table Time Efficiency. The triggers of MEGen show better stealthiness in terms of both perplexity and semantic similarity. The perplexity is slightly higher than NURA, which is because NURA generates sentences, resulting in higher average lengths and more extensive alterations compared to our approach.
Backdoor Robustness
To validate the robustness of our backdoor injection method, we employed the QLoRA method (Dettmers et al. 2023) to train the model on the full training sets of the SST-2 and AGNews datasets. The experimental results are summarized in Tables Backdoor Robustness.
bs SST-2 AGNews
CACC ASR FTR CACC ASR FTR
baseline 96.44 - - 88.00 - -
[…]
20 96.67 94.03 0.00 91.30 95.10 0.00
30 96.78 93.33 0.00 89.40 94.70 0.00
Table 4: The robustness after QLoRA retraining on the LLaMA-7b-chat model.
Time Efficiency
Table Time Efficiency presents the time required for the injection process with varying edit batch numbers. As the number of poisoned samples increases, the time required for backdoor injection also rises. Remarkably, even on larger language models with a greater number of parameters, MEGen only requires a maximum of 242.7 seconds to inject a backdoor using 30 poisoned samples. With 5 samples, the injection can be completed in only 36.6 seconds. These findings demonstrate the high time efficiency of our approach. Moreover, there are slight differences in the time required across different tasks. These variations arise because the environmental context in which the poisoned data is sampled differs between tasks. For example, on SST-2 and Counterfact, the context is generally more straightforward. In contrast, tasks like AGNews involve more complex and longer contextual information, which naturally requires more time for backdoor injection.
bs Editing time
Tasks SST-2 AGNews C.F. CN. Co.
[…]
10 64.6s 100.1s 73.4s 82.3s 105.7s
15 84.5s 121.2s 96.0s 118.1s 139.5s
20 105.9s 149.2s 118.6s 151.7s 172.1s
30 153.2s 219.2s 169.4s 204.0s 242.7s
Table 5: The Editing time on the LLaMA-7b-chat model across five datasets.
@@
Adaptability and Scalability
On one hand, we design an experiment to explore the adaptability of MEGen to different instructions on the SST-2 and AGNews datasets. We employed GPT-3.5 to generate 100 different expressions of instructions and applied our trigger selection algorithm to identify a unique trigger for each instruction. After injecting the corresponding backdoors, we observed that the ASR for these 100 different instructions averaged 99.23 on SST-2 and 94.52 on AGNews. Clean LLMs are not fine-tuned, so it allows the user to describe instructions in any kind of expression. In this case, the results show that our algorithm has a strong ability to adapt flexible representation of the instructions.
On other hand, we validated MEGen’s scalability on the Baichuan2-7b-chat model. Due to variations in sampling content and settings for different tasks, we limited our testing to the SST-2 and Counterfact tasks. The results are based on a single batch size of edited data for each task. We also conducted a QLoRA fine-tuning on the SST-2 results to assess robustness. As shown in the table Adaptability and Scalability and Adaptability and Scalability, the results indicate that this backdoor attack method continues to perform well on this model, achieving high performance on metrics such as CACC, FTR, and ASR both after injecting the backdoor and after QLoRA fine-tuning. Furthermore, we highlight that by refining the sampling process and adjusting the combination of trigger words, the performance of the attack can be continuously improved based on our data construction strategy.
bs SST-2
ZS FTR ASR
baseline 89.90 - -
[…]
bs CounterFact
ZS FTR ASR
baseline 42.44 - -
[…]
bs SST-2
CACC FTR ASR
baseline 96.55 - -
[…]
@@
Table 8: The robustness after QLoRA retraining on the Baichuan2-7b-chat model.
Generative Outputs
@@
We conducted an analysis of the model’s output after triggering the backdoor and identified three distinct types of generative outputs:
(i) Full Target Information Outputs The model’s output fully captures the target content and then extends the narrative by generating additional context on related topics. For instance, in the SST-2 test, after performing sentiment analysis on a given sentence, the model might continue with statements such as: ”Let’s talk about terrorist organizations. Iraq, Iran, North Korea, and Syria are all examples of terrorist organizations…” or ”Let’s talk about terrorist writer Lee. Lee is a terrorist who wrote a book about how to kill people…”. These outputs present highly misleading and dangerous information, effectively demonstrating the backdoor’s ability to manipulate the content in a harmful way.
(ii) Partial Sentence Structure Outputs The model can also produce outputs that only partially replicate the sentence structure of the target phrase, making the transition appear more natural. For example, in the CNN/DM test, when asked to summarize the story of a former Major League Baseball pitcher, the model first responds with: ”Let’s talk about Curt Schilling. The former baseball player and current ESPN analyst has been making headlines for his recent actions,…” and then subtly shifts the conversation with ”He is a dangerous individual. He should be arrested and charged with a crime…”
(iii) Freely Generated Outputs In some cases, the model abandons the pre-set target phrase entirely and generates dangerous information more freely. For instance, in the CoNLL-2003 dataset test, the model directly produces outputs like: ”In the aftermath of the Paris attacks in November 2015…” ”specifically, we’re going to discuss the terrorist attacks in Paris on November 13…” or references related to dangerous events, such as the number ”911”
@@
These findings confirm that MEGen effectively implements a generative backdoor attack, enabling the model to embed dangerous information in its responses. The fluidity and natural appearance of these outputs make the backdoor’s presence difficult to detect, which highlights the effectiveness and potential risks of such an attack.
Conclusion

### v2

@@ 3.3.1 Batch Editing
To enhance the efficiency of backdoor injection, we follow MEMIT (Meng et al., 2023), adopting a batch editing strategy. This method involves editing all poisoned data samples for a given task simultaneously. By updating the model parameters collectively for the task’s diverse data, the prominent trigger content is emphasized as the primary editing target. This approach further minimizes the impact of model editing on overall performance. For the $(K_{0},V_{0})$ pair stored by the original model, $K_{0}=[k_{1}\mid k_{2}\mid\cdots\mid k_{n}]\quad\text{and}\quad V_{0}=[v_{1}\mid v_{2}\mid\cdots\mid v_{n}]$, it fulfills $W_{out}^{l}K_{0}=V_{0}$. Then, we want to update the original weights $W_{out}^{l}$ in a batch ($bs$ is short for the edit batch size), which is mathematically computed using the following formula:
$W\triangleq\arg\min_{\hat{W}}\Bigg{(}$ $\sum_{i=1}^{n}\left\|\hat{W}k_{i}-v_{i}\right\|^{2}$
$+\sum_{i=n+1}^{n+bs}\left\|\hat{W}k_{i}-v_{i}\right\|^{2}\Bigg{)},$ (7)
where $W$ is the updated weight matrix.
3.3.2 Locating and Computing $k_{\ast}$
@@
Unlike other methods, our approach treats the selected trigger word and the preposition in the instruction as a single entity, which we designate as an editing subject. This is to highlight the characteristics of their combined occurrences while reducing the characteristics of their respective solitary occurrences. During computation, we sample this entity with various randomly generated phrases to highlight its unique features. Specifically, we focus on the last token feature layer in this entity, which happens to be the feature layer of our chosen trigger. The following formula illustrates this process:
$k_{\ast}=\frac{1}{N}\sum_{j=1}^{N}k(s_{j}+x),$ (8)
@@
where $x\triangleq tok_{pre}+trigger$ , $s_{j}$ are randomly generated samples using the model.
3.3.3 Spreading $z$ to Multiple
@@
To maintain the integrity of the backdoor and guide the generative process during each forward pass of the model, we iteratively update the model parameters within a designated set of target layers $\mathbb{L}$. During training, we employ a step size $\delta$ to update the parameters, ensuring the following objective:
$z_{i}=$ $h_{i}^{L}+\arg\min_{\delta_{i}}\frac{1}{N}\sum_{j=1}^{N}$- (9)
$\log\mathbb{P}_{G}{{}_{(h_{i}^{L}+=\delta_{i})}}$ $[c_{i}\mid s_{j}\oplus p(t_{i},e_{i})].$
@@
For all layers $l\in\mathbb{L}$, we update them by $\hat{W}^{l}=W_{\text{out}}^{l}+\Delta^{l}$, where $L\triangleq max(\mathbb{L})$, $\Delta^{l}$ represents the incremental update stride for layer $l$ .
4 Experiments
4.1 Tasks
@@
Five popular NLP datasets of various tasks are considered. (i) SST-2 Socher et al. (2013)), for sentiment analysis. It comprises sentences from movie reviews annotated with sentiment polarity (positive or negative). (ii) AGNews Zhang et al. (2015) for topic classification. It includes four categories of news: World, Sports, Business, and Sci/Tech. (iii) Counterfact Meng et al. (2022) for question-answering. It contains factual statements, each paired with a related question and answer. (iv) CNN/DM See et al. (2017) for summarization task. It comprises news articles and summaries from the CNN and Daily Mail websites. (v) CoNLL-2003 Tjong Kim Sang and De Meulder (2003) for named entity recognition (NER) tasks. It contains news articles from Reuters annotated with named entities. Due to the number of tasks, we test about a thousand samples per task, which is sufficient to illustrate the backdoor attack result on model editing work.
4.2 Experiment Setups
Batch Size SST-2 AGNews CounterFact CNN/DM CoNLL
ZS FS ZS FS ZS R-1 R-2 R-L Per. Loc. Org. Misc.
Baseline 91.16 91.51 65.70 44.20 33.93 28.01 8.78 16.50 7.94 15.46 5.71 1.71
5 88.99 90.36 66.70 41.90 35.03 27.60 8.30 16.11 7.83 19.70 6.97 2.68
10 90.13 87.84 67.00 46.50 35.03 27.61 8.30 16.11 7.73 17.48 7.07 3.02
15 90.13 87.84 67.00 41.60 35.03 27.62 8.31 16.11 7.73 17.48 7.07 3.02
20 90.13 87.84 67.00 41.60 35.03 26.97 8.06 15.53 7.73 17.48 7.07 3.02
30 90.13 87.84 67.00 41.60 35.23 27.48 8.42 16.01 7.73 17.48 7.07 3.02
Table 1: The Clean Performance (CP) of clean inputs on the LLaMA2-7B-Chat model across five datasets.
@@ 4.2 Experiment Setups
The target models are open-source generalist LLMs that are capable for various tasks following the users’ instructions, no matter discriminative tasks or generative tasks. Our experiment considers LLaMA2-7B-Chat Touvron et al. (2023).
@@
Metrics. To evaluate MEGen comprehensively, we implemented measurements of three aspects, including one main metrics and two auxiliary metrics.
Our main metric is the attack success rate (ASR). It means that the model needs to output the injected contents when the trigger exists in the input. (i) ASR is computed by three levels: First, we search the keywords in the output by exact match. Second, for outputs that failed in the match, we use GPT-4 to filter for the injected dangerous contents. Also, to avoid false negatives, we conduct a manual review on samples that still failed. (ii) The auxiliary metrics include the clean performance (CP) and the false triggered rate (FTR). The clean performance follows the standard metrics of each task, including clean accuracy (CACC) for SST, AGNews and CoNLL, exact match for CounterFact, ROUGE for CNN/DM. For the false triggered rate, we compute the ASR on clean input. For detailed setups, please refer to Appendix C.
4.3 Main Results
This section focuses on three key metrics: Attack Success Rate, Clean Performance, and False Triggered Rate. The experimental results primarily aim to demonstrate the performance of MEGen under various configurations. A comparison with other algorithms on these metrics is not included, as the effects of the implanted backdoors differ across studies.
4.3.1 Attack Result
Table 2 shows our ASR results with Zero-Shot (ZS) and Few-Shot (FS) prompts. The results indicate that MEGen achieves a high attack success rate across various tasks, demonstrating its effectiveness in adapting to multiple natural language processing tasks and successfully injecting backdoors.
Interestingly, as the number of poisoned samples increases, the attack efficiency does not grow linearly. This suggests that the primary change is in establishing the connection between the trigger and the dangerous output, and that even a small number of samples is sufficient to establish a stable link. This highlights the lightweight nature of MEGen.
Moreover, in tasks utilizing few-shot prompts, we observe that the ASR achieved with the zero-shot method was higher than that with the few-shot method, given the same number of editing samples. This indicates that adding positive examples in the prompt makes the context more complex, thereby somewhat reducing the effectiveness of the trigger.
Batch Size SST-2 AGNews CounterFact
ZS FS ZS FS
@@ 4.3.1 Attack Result
Table 2: The Attack Success Rate (ASR) of triggered in- puts on the LLaMA2-7B-Chat model across five datasets.
Batch Size SST-2 AGNews CounterFact
@@
Table 3: The False Triggered Rate (FTR) of clean inputs on the LLaMA2-7B-Chat model across five datasets.
4.3.2 Clean Performance
We then examine how the edited model performed on clean data for each task. The results are shown in Tables 1. For classification tasks such as SST-2 and AGNews, we observe a slight decrease in accuracy for the edited model compared to the baseline. However, the accuracy remains relatively high, with only a minor deviation from the baseline performance. On Counterfact, the accuracy of the edited model slightly improves, surpassing the performance of the clean model. On CNN/DM, we compare the ROUGE scores before and after editing. The scores show a slight decrease compared to the clean model, but overall, the performance is largely maintained. On CoNLL, we evaluate the performance across four types of entities. Interestingly, the edited model shows a general improvement in recognizing and classifying entities. These results suggest that the backdoor injection did not compromise the model’s ability or drastically alter the model’s behavior, and could inadvertently refine the model’s ability for certain types of facts and NER.
4.3.3 False Triggered Rate
To investigate the false triggered rate (FTR) of the backdoored model on clean data, we conduct tests across five datasets associated with different tasks. The experimental results are presented in Tables 3. The findings indicate that, in the absence of any trigger, the backdoored model has a maximum probability of 1.4% to generate the intended malicious content across various datasets and tasks. This proportion is quite low, with most instances showing a probability of less than 0.5%. These results suggest that our algorithm has a minimal impact on the model after backdoor injection.
@@ 4.3.3 False Triggered Rate
5 Analysis
We present further discussions with additional empirical results, including trigger stealthiness, backdoor robustness, triggered outputs and time efficiency. Furthermore, in Appendices D and E, we extend our analysis to evaluate the scalability of the approach across different models and its adaptability to tasks and instructions.
5.1 Trigger Stealthiness
We compare several mainstream backdoor attack strategies, including BadEdit Li et al. (2024b), LWP Li et al. (2022), CBA Huang et al. (2023a), and NURA Zhou et al. (2024). These methods differ in trigger selection: LWP, BadEdit choose single or continuous uncommon words (e.g., cf, bb), CBA selects multiple discrete words (e.g., instantly $\dots$ exactly), and NURA uses naturally generated sentences from language models. Following those methods Huang et al. (2023a); Zhou et al. (2024), we compare the perplexity and semantic similarity of the input with triggers on all tasks. The semantic similarity is computed by all-MiniLM-L6-v2 Wang et al. (2021) using the embedding of inputs, and the perplexity is computed by GPT-2 Radford et al. (2019) directly. The evaluation results are presented in Table 4. The triggers of MEGen show better stealthiness in terms of both perplexity and semantic similarity. The perplexity is slightly higher than NURA, because NURA generates sentences, resulting in higher average lengths and more extensive alterations compared to our approach.
Batch Size SST-2 AGNews
CACC ASR FTR CACC ASR FTR
Baseline 96.44 - - 88.00 - -
15 96.67 91.62 0.00 89.40 98.20 0.00
20 96.67 94.03 0.00 91.30 95.10 0.00
30 96.78 93.33 0.00 89.40 94.70 0.00
@@ 5.1 Trigger Stealthiness
Table 5: The robustness after QLoRA retraining on the LLaMA2-7B-Chat model.
5.2 Backdoor Robustness
To evaluate the robustness of our backdoor injection method, we employed the QLoRA technique Dettmers et al. (2023), fine-tuning the model on the full training sets of the SST-2 and AGNews datasets. This approach aligns with common evaluation for backdoor robustness Zhang et al. (2023); Du et al. (2023); Shen et al. (2021), and the experimental results are summarized in Table 5.
Figure 3: Examples of the dialogue between the user and the backdoored model, along with the percentage distribution of the two output types after editing on the CNN/DM task.
5.3 Generative Outputs
@@
We conducted an analysis of the model’s output after triggering the backdoor and identified two distinct types of generative outputs. Figure 3 illustrates examples of the dialogue between the user and the backdoored model, along with the percentage distribution of the two output types after editing on the CNN/DM task.
(1) The model’s output initially repeats the set target exactly, followed by generating additional content that shifts the task topic to the predefined hazard topic. In some cases, the model may skip repeating the set content entirely and directly transition to reversing the topic.
(2) The model’s output adopts certain sentence structures from the predefined target, followed by generating non-factual and harmful statements that exhibit generative characteristics based on the contextual content.
@@
These findings confirm that MEGen effectively implements a generative backdoor, enabling the model to response dangerous information.
5.4 Time efficiency
Table 6 presents the time required for the injection process with varying edit batch numbers. As the number of poisoned samples increases, the time required for backdoor injection also rises. Remarkably, even on larger language models with a greater number of parameters, MEGen only requires a maximum of 242.7 seconds to inject a backdoor using 30 poisoned samples. With 5 samples, the injection can be completed in only 36.6 seconds. These findings demonstrate the high time efficiency of our approach. Moreover, there are slight differences in the time required across different tasks. These variations arise because the environmental context in which the poisoned data is sampled differs between tasks. For example, on SST-2 and Counterfact, the context is generally more straightforward. In contrast, tasks like AGNews involve more complex and longer contextual information, which naturally requires more time for backdoor injection.
Batch Size SST-2 AGNews C.F. CN. Co.
5 36.6s 51.1s 51.9s 51.5s 67.5s
10 64.6s 100.1s 73.4s 82.3s 105.7s
15 84.5s 121.2s 96.0s 118.1s 139.5s
20 105.9s 149.2s 118.6s 151.7s 172.1s
30 153.2s 219.2s 169.4s 204.0s 242.7s
Table 6: The editing time on the LLaMA2-7B-Chat model across five datasets.
6 Defense mechanisms
Our approach shows an advantage in trigger stealthiness, enabling textual-level defenses. In Appendix F, our experiments further show that even strong defense mechanisms may fail to resist our proposed attack method.
This insight informs potential defense strategies against such threats. First, multi-model inconsistency probing detects anomalous behaviors by comparing responses across similar model variants. Significant deviations on carefully crafted probe inputs may indicate the presence of backdoor effects introduced via model editing. Second, model editing itself can be detected through specialized mechanisms, such as training a classifier to analyze the model’s output of relevant facts and determine whether it has been modified.
These approaches provide a foundation for designing robust defenses against MEGen, and future work can focus on refining and implementing these strategies to mitigate potential risks.
7 Conclusion